RC Meeting and Run Plan, for Friday, March 6, 2020
- 1 RC Meeting Information
- 2 General updates:
- 3 Current Run Plan:
- 4 Bluejeans remote connection information
- 5 Useful Links
RC Meeting Information
RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection
Friday March 6, 2020.
- No more issues with SHMS dipole PS trips. It seems the replacement of the timer module (used in a detect-heartbeat-or-shut-the-hell-down application) did the trick.
- Moeller run was done on Fri day. (85.2+-0.4%) Thanks, guys! Opportunistic pNMR tests as well.
- Took HMS/SHMS -3.5/-3.4 GeV data over Thursday nite with 3He target, longitudinal polarization. Good production.
- Three target polarization measurements since restoration yesterday: 51%, 57%, and 53%.
- Beam delivery Friday day problematic due to a serious trip in cryomodule 1L23, and then an injector issue. Longitudinal production resumed at 17:25 Friday.
- Rapid progress was made Thursday nite in (mostly) resolving the target wall scraping issue that Bill Henry noted. Reducing the FR size from 4.5mm diameter to 4.0mm seemed to improve things, suggesting it was real scraping and not a FR artifact. https://logbooks.jlab.org/entry/3798113 Then JP shifted the beam position by 0.5mm in the correct direction(!) and it got even better, proving that scraping was involved. (confusing assortment of hclog entries on this) With 20/20 hindsight, it had been noticed that the burn scar on this cell window was offset, and the original carbon hole target plot indeed had a 1mm offset. https://logbooks.jlab.org/entry/3782272 So the apparatus is reproducible, but the sensitivity to beam position seems surprisingly high https://logbooks.jlab.org/entry/3798550 . Simona suggested checking the YAG viewer which showed no obvious nasty tail. https://logbooks.jlab.org/entry/3798716 Xiaochao et al will study the history.
Current Run Plan:
Friday DAY thru Saturday OWL
- COMPLETE! Moller measurement during day shift beginning 08:30. When Moller crew is ready, end production and take one NMR measurement, then let Moller crew take over.
- Once Moller is completed:
- Follow the Short beam recovery procedure(Skip the YAG unless you are concerned about the harp scan data.)
- Do one NMR before resuming production
- Resume production with the 3He target in longitudinal direction, (HMS/SHMS -3.5/-3.4 GeV).
- Then one NMR every 5 hours.
- Sometime in late owl/early Day shift, after one scheduled NMR measurement (in longitudinal), rotate target to transverse (90 deg, rotation by expert only).
- Take 5 hours of transverse production data there.
- Upon completion, do one NMR in the transverse direction, then rotate target to longitudinal (180 deg, rotation by expert only).
- Resume longitudinal production data taking with target at 180 deg, take NMR every 5 hours.
Sunday and beyond: ideally would go like Saturday plan. We'll see.
Notes for the Future
- 'Filler' runs for next spin up / target-down period (Don't want to forget about this!):
- 1 hour run with target out for measure of the background just from the beryllium windows. (Needs SHMS, btw.)
- pNMR issues
- Software/driver issue needs to be resolved
- NO HCLOG ENTRY Checkout during next non-production down time (Moller or SHMS fix)
- EPR problems  saw some progress on Tuesday , but issues remain
- Tentative plan now is to continue the d2n (and A1n) program by replacing the photo-diode that is placed directly on the pumping chamber (we have many spares)
- It would be good to write up a procedure to streamline this replacement procedure for us and RadCon
- Kepco power supply commissioning/installation (Todd, Bill)
- Bill will update status on Thursday
- Verify that calibration constants and input values associated with NMR and other polarimetry are consistent and will be correct when the software is restarted      (Bill, Junhao)
- Persistent BLM & ion chamber trips
- General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
- NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
- Beam Studies program will investigate some of these issues (most pressing is significant activation + beam loss in the arcs)
- Contact with Target readbacks through AnywhereUSB USB bridge still drop out on occasion
- Initial replacement of the USB bridge failed (did not seem to help underlying issue, then bridge stopped working), replaced with original device
- OPs wished to reboot hung iochc10 , noted in December as well 
- This runs the Moller cryo controls and should not be rebooted without expert oversight (does not impact operations). D. Gaskell is looking into getting this addressed properly with OPs.
Opportunistic Access Jobs
None at the moment.
If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change the beam tuning, then do [https://hallcweb.jlab.or/wiki/index.php/Beam_and_Raster_Checkout_for_A1n/d2n_(Short_Procedure) beam recovery procedure] (do YAG viewer part only if you are suspicious of the harp scan results)
- NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell the Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
- Check the position locks are on.
- Do your best to accommodate Hall A's request for a PITA scan promptly. It is OK to stop the present runs and start new ones in 5--10 minutes after they are done.
- Check that CODA is running normally (had one strange hangup lately)
- Check online replays shortly after starting runs, & full replay results too.
- Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
- Always take 120Hz of EDTM
- Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! !! !! !! ATTN: SHIFT LEADERS !! !! Please follow the shift summary template on !! !! the laminated SL instructions. !! !! !! !! ATTN: TARGET OPERATORS !! !! All NMR measurement HCLOG entries should include: !! !! 1) A title that clearly lists polarization !! !! 2) Screenshots of both NMR_FieldSweep and !! !! the polarization fit !! !! !! !! For an example see the following entry: !! !! https://logbooks.jlab.org/entry/3768696 !! !! !! !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
Notices in Effect for Shift Crew
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
- Spectrometer GUI (
- Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
- Ramp rate lock instructions are here and added to checklist
Bluejeans remote connection information
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-184.108.40.206 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920
Target coil rotation planned for March 13--20
- Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program
- More details are here: PolHe3_Target_Coil_Rotation_Plan,_Mar_2020
Target Polarization Angle Conventions
Be VERY clear here about the convention (beam-right/left/up/downstream) vs 0, 90, 180, 270) for each run period. More notes and references to log entries are better.
Polarization Calibration Data
ADD LIST OF CALIBRATION RUNS HERE AND A TABLE OF RELEVANT CALIBRATION CONSTANTS TO USE FOR WHAT RUNS
| SHMS Angle
| SHMS Momentum
| Field Direction
|Calibration Results||Status|| Oven Temperature
|Tommy||EPR-NMR||11||7.5||90||5.24 %/mV||Done||200|| 3839939|
|Tommy||AFP Loss||14.5||6.4||90||1.25%/Sweep||Done||200||3840618, 3840143, 3840618|
|Tommy|| EPR Amplifier
on/off NMR Signal
Target Polarization Database
d2n Run Period
- Junhao Chen is maintaining it
A1n Run Period
- Mingyu Chen will updating the runlist while Murchhana Roy will do the cross check.
Recent polarization measurements
- longitudinal polarization measurements after yesterday's target work in the hall and QWP angle optimization
- Tue 2/18 13:15: 47% https://logbooks.jlab.org/entry/3786686
- Tue 2/18 18:42: 52.2% https://logbooks.jlab.org/entry/3786898
- Tue 2/18 23:47: 54.5% https://logbooks.jlab.org/entry/3787175
- Wed 2/19 4:38: 55.0% https://logbooks.jlab.org/entry/3787175
- Wed 2/19 8:40: 54.9% https://logbooks.jlab.org/entry/3787285
- Wed 2/19 13:08 54.8% https://logbooks.jlab.org/entry/3787420
- Wed 2/19 18:15 54.8% https://logbooks.jlab.org/entry/3787605
- Wed 2/19 11 PM 55.3% https://logbooks.jlab.org/entry/3787781
- Wed 2/19 midnight 51.9% https://logbooks.jlab.org/entry/3787926
- Thu 2/20 05:15 53.2% https://logbooks.jlab.org/entry/3787904
- Thu 2/20 10:34 53.6% https://logbooks.jlab.org/entry/3788085
- Thu 2/20 13:47 53.0% https://logbooks.jlab.org/entry/3788210
- Thu 2/20 21:00 56.0% https://logbooks.jlab.org/entry/3788428
- Fri 2/21 04:07 55.2% L https://logbooks.jlab.org/entry/3788613
- Fri 2/21 09:03 53.8% L https://logbooks.jlab.org/entry/3788795
- Fri 2/21 14:22 52.5% L https://logbooks.jlab.org/entry/3788938
- Fri 2/21 19:50 51.9% L https://logbooks.jlab.org/entry/3789160
- Sat 2/22 01:15 51.6% L https://logbooks.jlab.org/entry/3789331 -> QWP optimization done
- Sat 2/22 06:32 53.0% L https://logbooks.jlab.org/entry/3789442
- Sat 2/22 11:50 54.9% L https://logbooks.jlab.org/entry/3789583 -> rotated to T (5hr) then back to L (2-3hrs)
- Sat 2/22 19:31 43.2% L https://logbooks.jlab.org/entry/3789857
- Sat 2/22 23:08 50.2% L https://logbooks.jlab.org/entry/3789959
- Sun 2/23 04:06 55.3% L https://logbooks.jlab.org/entry/3790104
- Sun 2/23 09:21 55.5% L https://logbooks.jlab.org/entry/3790240
- Sun 2/23 14:17 55.7% L https://logbooks.jlab.org/entry/3790380
- Sun 2/23 19:03 55.7% L https://logbooks.jlab.org/entry/3790513
- Mon 2/24 00:25 55.2% L https://logbooks.jlab.org/entry/3790672
- Mon 2/24 05:47 49.6% L https://logbooks.jlab.org/entry/3790771 -> unexplained drop
- Mon 2/24 10:59 50.8% L https://logbooks.jlab.org/entry/3790965
- Mon 2/24 15:51 50.4% L https://logbooks.jlab.org/entry/3791100
- Mon 2/24 21:22 50.9% L https://logbooks.jlab.org/entry/3791318
- Tue 2/25 02:59 50.7% L https://logbooks.jlab.org/entry/3791481 -> QWP optimization done
- Tue 2/25 08:30 51.4% L https://logbooks.jlab.org/entry/3791660
- Tue 2/25 18:13 ~50% L https://logbooks.jlab.org/entry/3791950
- Tue 2/25 23:13 53.3% L https://logbooks.jlab.org/entry/3792124
- Wed 2/26 04:34 53.6% L https://logbooks.jlab.org/entry/3792291
- Wed 2/26 09:04 54.3% L https://logbooks.jlab.org/entry/3792418 -- NMR just after beam stops
- Wed 2/26 12:46 50.1% L https://logbooks.jlab.org/entry/3792504 -- NMR before beam returns
- Wed 2/26 18:44 52.8% L https://logbooks.jlab.org/entry/3792704
- Wed 2/26 23:42 53.8% L https://logbooks.jlab.org/entry/3792849
- Thu 2/27 04:43 56.4% L https://logbooks.jlab.org/entry/3792931 -- not much beam during owl
- Thu 2/27 08:38 56.3% L https://logbooks.jlab.org/entry/3793085
- Thu 2/27 08:38 56.3% L https://logbooks.jlab.org/entry/3793085
- Thu 2/27 16:50 58.4% L https://logbooks.jlab.org/entry/3793359 -- not much beam during day shift (Moller + retune)
- Thu 2/27 21:49 58.0% L https://logbooks.jlab.org/entry/3793493
- Fri 2/28 02:53 57.8% L https://logbooks.jlab.org/entry/3793610
- Fri 2/28 07:41 57.9% L https://logbooks.jlab.org/entry/3793727
- Fri 2/28 12:54 58.5% L https://logbooks.jlab.org/entry/3793881 -- just before beam restart; after EPR hardware misconfiguration was corrected
- Fri 2/28 18:04 57.2% L https://logbooks.jlab.org/entry/3794035
- Fri 2/28 18:19 49.8% (55.7%) T https://logbooks.jlab.org/entry/3794045 -- value in () is with updated calib. coefficient
- Fri 2/28 20:21 49.2% (55.1%) T https://logbooks.jlab.org/entry/3794079 -- value in () is with updated calib. coefficient
- Sat 2/29 00:37 --- % T https://logbooks.jlab.org/entry/3794242 -- Transverse NMR did not work
- Sat 2/29 01:03 52.5% L https://logbooks.jlab.org/entry/3794244
- Sat 2/29 06:10 54.3% L https://logbooks.jlab.org/entry/3794356
- Some problematic NMR measurements around 12:00--12:14? 
- Sat 2/29 12:25 52.8% L https://logbooks.jlab.org/entry/3794512
- Sat 2/29 17:49 54.8% L https://logbooks.jlab.org/entry/3794693
- Sat 2/29 22:51 54.4% L https://logbooks.jlab.org/entry/3794802
- Sun 3/01 03:58 53.9% L https://logbooks.jlab.org/entry/3794944
- Sun 3/01 09:42 54 % L https://logbooks.jlab.org/entry/3795082
- Sun 3/01 11:00 ----- https://logbooks.jlab.org/entry/3795123 -- Re-analyzed some odd sweeps from Sat DAY shift
- Sun 3/01 14:33 54.5% L https://logbooks.jlab.org/entry/3795265
- Sun 3/01 19:58 54.1% L https://logbooks.jlab.org/entry/3795414
- Mon 3/02 01:06 54.1% L https://logbooks.jlab.org/entry/3795560
- Mon 3/02 06:08 53.9% L https://logbooks.jlab.org/entry/3795662
- Mon 3/02 11:08 53.9% L https://logbooks.jlab.org/entry/3795836 -- before switching to Transverse
- Mon 3/02 11:32 52.5% T https://logbooks.jlab.org/entry/3795851
- Mon 3/02 15:04 54.4% T https://logbooks.jlab.org/entry/3795982 -- just before production beam
- Mon 3/02 17:41 https://logbooks.jlab.org/entry/3796117 -- NMR PC Transverse Calibration constant updated
- Mon 3/02 20:37 53.9% T https://logbooks.jlab.org/entry/3796180
- Mon 3/02 20:46 52.9% L https://logbooks.jlab.org/entry/3796182
- Tue 3/03 00:47 55.6% L https://logbooks.jlab.org/entry/3796277
- Tue 3/03 05:36 54.7% L https://logbooks.jlab.org/entry/3796415
- Tue 3/03 08:19 54.7% L https://logbooks.jlab.org/entry/3796485